OpenStudy (anonymous):

2x+y=3 3x-y=2 write the solution in the form (x,y)

OpenStudy (a1234):

2x+y=3 3x-y=2 I''ll use elimination: 5x = 5 x = 1 2(1) + y = 3 2 + y = 3 y = 1 (1, 1)